OpenStudy (anonymous):

six burned-out bulbs have been mixed in with 16 good ones. Ken is replacing old bulbs in his house. If he selects two bulbs at random from the box of 22, what is the probability they both work? P(Both work) = 64/121 P(Both work) = 61/121 P(Both work) = 8/11 P(Both work) = 40/77

OpenStudy (welshfella):

probability of the first one picked being OK = 16/22 or 8/11 what do you think probability of 2nd one being ok is?

OpenStudy (anonymous):

15/21 ??

OpenStudy (welshfella):

exactly right which reduces to 5/7 right?

OpenStudy (anonymous):

yes, but why isnt it one of the possible answers that is why i am confused

OpenStudy (anonymous):

oh wait, i have to multiply them now right

OpenStudy (welshfella):

because you have 2 events so you have either to add or multiply the probabilities. In this case we multiply them because the 2 'picks' are independent of each other. so final answer is 3/11 * 5/7

OpenStudy (anonymous):

i see

OpenStudy (welshfella):

good

OpenStudy (anonymous):

8/11 x 5/7 = 40/77 answer is D

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

correct a little more than 50% yw